# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

### v4

@@ 5.2 Qualitative results
Table 1: Segmentation (IoU) on NVOS (Ren et al., 2022) and SPIn-NeRF (Mirzaei et al., 2023).
Geometry only Single view Uplifting Uplifting + Graph diffusion
@@
Table 2: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023). We compare purely geometrical reference mask uplifting and reprojection and single-view prediction, feature/mask uplifting or graph diffusion leveraging DINOv2 or SAM2.
5.3 Segmentation results
In this section, we quantitatively evaluate the segmentation task on NVOS (Ren et al., 2022) and SPin-NeRF (Mirzaei et al., 2023). We evaluate segmentation based on SAM and SAM2 mask uplifting, and on DINOv2 feature uplifting combined with graph diffusion. We compare our segmentation results to the current state of the art: SA3D (Cen et al., 2023c), SA3D-GS (Cen et al., 2023b), SAGA (Cen et al., 2023a), OmniSeg3D (Ying et al., 2024). All these methods are specifically designed for uplifting the 2D segmentation masks produced by SAM into 3D using gradient-based optimization of a projection loss. We also report results from NVOS (Ren et al., 2022) and MVSeg (Yen-Chen et al., 2022), who respectively introduced the NVOS and SPIn-NeRF datasets.
Results.
Table 1 reports the average IoU across all scenes for NVOS and SPIn-NeRF. Per-scene results can be found in Appendix Tables 6 and  7. Our results are comparable to the state-of-the-art, while not relying on gradient-based optimization. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Compared to SAM, DINOv2 better captures complex objects, but sometimes also captures some background noise. This can be seen in Appendix Figure 5 with the example of the T-Rex: while SAM misses out the end of the tail as well as the end of the ribs, DINOv2 captures the whole Trex, but also captures part of the stairs behind. Our lower segmentation results compared to OmniSeg’s can partly be attributed to the poor Gaussian Splatting reconstruction of highly specular scenes, such as the Fork. As also noted by Cen et al. (2023a), the reconstruction includes semi-transparent Gaussians floating over the object, attempting to represent reflections or surface effects, which are challenging to capture using standard rasterization techniques (Jiang et al., 2024).
Ablation study.
We compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate i) a purely geometrical variant that reprojects the reference mask on the other views, without using SAM2 or DINOv2, ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, and iv) segmenting using graph diffusion over DINOv2 3D feature similarities. Results are reported in Table 2, and per-scene IoU as well as a detailed analysis can be found in Appendix Table 8 and Sec. B.2. We observe that the purely geometrical approach works well on the forward-facing scenes and fails on 360-degree scenes. The single-view variant performs reasonably well on average but, the low resolution of patch-level representations (illustrated in Figure 2) lead to a coarser segmentation. 3D uplifting considerably boosts results compared to single-view approaches, and introducing 3D spatial information through 3D graph diffusion further enhances results on the more challenging 360-degree scenes.
5.4 Open-vocabulary object detection
Table 3 presents results on the open-vocabulary LERF localization and segmentation tasks (Kerr et al., 2023), evaluated on the extended and more challenging version of the dataset introduced by LangSplat (Qin et al., 2024). The reported average running times include feature map generation and 3D feature training whenever relevant. LUDVIG outperforms prior works while being significantly faster than all methods (around 10 times faster). Table 3 reports our results for segmentation using graph diffusion for refining 3D CLIP relevancy scores, followed by segmentation with SAM. Table 4 also reports results with and without using graph diffusion and/or SAM.
A more thorough analysis on running times can be found in Appendix B.1. Additionally, Appendix C.4 provides illustrations of the impact of the 3D diffusion process and 2D SAM segmentation (Figure 10), and comparative visualizations of localization heatmaps with LangSplat and LERF (Figure 11).
Localization (Accuracy) Segmentation (IoU) Time
ramen figurines teatime waldo overall ramen figurines teatime waldo overall (mins)
LERF 62.0 75.0 84.8 72.7 73.6 28.2 38.6 45.0 37.9 37.4 45
LangSplat 73.2 80.4 88.1 95.5 84.3 51.2 44.7 65.1 44.5 51.4 105
LUDVIG 78.9 80.4 94.9 90.9 86.3 58.1 63.3 77.1 58.5 64.3 10
@@ 5.4 Open-vocabulary object detection
Table 3: LERF Localization and Segmentation. We evaluate on the more challenging dataset introduced by LangSplat (Qin et al., 2024).
@@
Table 4: LERF Segmentation. Results (IoU) with and without using 3D graph diffusion and/or 2D SAM segmentation, evaluated on the dataset introduced by LangSplat (Qin et al., 2024).
@@ 6 Concluding remarks and limitations
In this work, we introduce a simple yet effective aggregation mechanism for transferring 2D visual representations into 3D, bypassing traditional optimization-based approach. The aggregation builds upon already trained Gaussian Splatting representations and is implemented within the CUDA rendering process, making 2D-to-3D uplifting as fast as 3D-to-2D rendering. Note however that the quality of 3D features is bound by that of the 3D scene reconstruction. Reconstruction by Gaussian Splatting is notoriously challenging in cases of high specularity (Jiang et al., 2024; Yang et al., 2024), blurred images (Zhao et al., 2024; Lee et al., 2024) or high-frequency regions (Ye et al., 2024b; Zhang et al., 2024). In such scenarios, learning 3D features along with 3D Gaussian Splatting reconstruction may improve scene geometry, opening promising perspectives for future work.
After feature uplifting, our graph diffusion process allows injecting the rich DINOv2 representations to transform coarse 2D or 3D segmentation signals, such as scribbles or CLIP relevancy scores, into accurate 3D segmentation masks. Our CLIP relevancy refinement builds upon prior works using DINO features as a regularization (Kerr et al., 2023; Zuo et al., 2024), while alleviating the computational overhead associated with joint gradient-based optimization of CLIP and DINO features.
@@ Appendix A Using LUDVIG for downstream tasks
In this section, we describe our approach for uplifting DINOv2, SAM and CLIP models and evaluating the 3D features on two downstream tasks: segmentation and open-vocabulary object detection.
As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$ and a corresponding 3D scene obtained using the Gaussian Splatting method.
Multiple-view segmentation.
For this task, we assume that a foreground mask of the object to be segmented is provided on a reference frame taken to be the first frame $I_{1}$. We consider two types of foreground masks: either scribbles or a whole reference mask of the object, both of which define a set of foreground pixels $\mathcal{P}$. In the following subsections, we present the proposed approaches for segmentation using SAM and DINOv2 features, based on both types of foreground masks.
@@ A.1 Multiple-view Segmentation with SAM
SAM (Kirillov et al., 2023; Ravi et al., 2024) is a powerful image segmentation model, that can generate object segmentation masks from point prompts on a single 2D image. Aggregating SAM 2D segmentation masks in 3D allows for cross-view consistency and improves single-view segmentation results. In order to leverage SAM, we propose a simple mechanism for generating SAM 2D features for each frame from a foreground mask in the reference frame.
Generating 2D query points for SAM.
The key idea is to generate point prompts on each training frame from the foreground mask provided on the reference frame. To this end, we perform an uplifting of the foreground mask (Eq. (3)) and re-project it on all frames (Eq. (4)). From the reprojected mask for viewing direction $d$, further normalized by its average value, we retain a subset $\mathcal{P}_{d}$ of pixels with values higher than a threshold $\tau$ fixed for all scenes ($\tau=0.4$ for SPIn-NeRF and $\tau=1$ for NVOS). We then predict a SAM mask based on these point prompts as described below.
Predicting a SAM segmentation mask from a set of query points.
Given a set of pixels $\mathcal{P}_{d}$ pertaining to the foreground, we compute 2D segmentation masks using SAM by randomly selecting $3$ points prompts from $\mathcal{P}_{d}$, repeating the operation $10$ times and averaging the resulting masks for each view to obtain the final 2D SAM feature maps.
Segmentation with uplifted SAM masks.
The 2D segmentation masks generated by SAM are uplifted using the aggregation scheme described in Sec. 3.2. Our final prediction is obtained by rendering the uplifted feature maps into the target frame and thresholding.
Evaluation and hyperparameter tuning.
Segmentation with 3D SAM masks requires setting a threshold for foreground/background pixel assignment, and optionally choosing one of the three masks proposed by SAM (representing different possible segmentations of the object of interest). For SPIn-NeRF, the threshold and mask prediction are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided; hence, a single mask is predicted, and the segmentation threshold is fixed across all scenes for SAM, and automatically chosen using Li’s iterative Minimum Cross Entropy method (Li & Lee, 1993) for SAM 2.
@@ A.2 Multiple-view segmentation with DINOv2
DINOv2 (Oquab et al., 2024) is a self-supervised vision model recognized for its generalization capabilities. In this work, we aggregate the patch-level representations produced by DINOv2 with registers (Darcet et al., 2024) into a high resolution and fine-grained 3D semantic representation.
Construction of 2D feature maps.
We construct the 2D feature maps using a combination of a sliding windows mechanism and dimensionality reduction of the original DINOv2 features. Specifically, we i) extract DINOv2 patch-level representations across multiple overlapping crops of the training images, ii) apply dimensionality reduction over the set of all patch embeddings, ii) upsample and aggregate the dimensionality-reduced patch embeddings to obtain pixel-level features for each image. The process is illustrated in Figure 4. This approach enhances the granularity of spatial representations by aggregating patch-level representations to form pixel-level features. To favor the first principal components, known to focus on the foreground objects (Oquab et al., 2024), the features are re-weighted by the eigenvalues of the PCA decomposition.
Segmentation with uplifted DINOv2 features.
The 2D feature maps from the $m$ training views are uplifted using Eq. (3) and the resulting 3D features are then re-projected into any viewing direction $d$ using Eq. (4) to compute rendered 2D features $(\hat{F}_{d,p})$. To obtain segmentation masks, we construct a score $P(\hat{F}_{d,p})$ for a 2D pixel $p$ to belong to the foreground, based on its corresponding rendered feature. More precisely, $P$ relies on the rendered foreground features $\mathcal{F}_{\text{ref}}:=(\hat{F}_{d_{1},p})_{p\in\mathcal{P}}$ corresponding to the foreground mask computed on the reference frame $I_{1}$. We propose two approaches for constructing $P$. The first one is a simple approach that sets $P(\hat{F}_{d,p})=\mathcal{S}_{F}(\hat{F}_{d,p},\bar{F})$ where $\bar{F}$ is the average over foreground features $\mathcal{F}_{\text{ref}}$, and $\mathcal{S}_{F}$ is defined based on the cosine similarity. The second approach is more discriminative and first trains a logistic regression model $P$ on all rendered 2D features of the reference frame, so that the foreground features $\mathcal{F}_{\text{ref}}$ are assigned a positive label. Then $P(\hat{F}_{d,p})$ gives the probability that a pixel $p$ belongs to the foreground. The final mask is then obtained by thresholding.
@@
Figure 4: Sliding windows for construction of DINOv2 feature maps.
@@ A.3 Enhancing segmentation with DINOv2 using 3D graph diffusion
DINOv2 provides generic visual features that do not explicitly include information for segmentation, unlike models such as SAM that were specifically trained for such a task. Consequently, using the 2D projections of uplifted DINOv2 features, as proposed in Sec. A.2, might fail to separate different objects that happen to have similar features while still being distinct entities. This challenge can be mitigated by incorporating 3D spatial information in which the objects are more likely to be well-separated. To this end, we propose to leverage the graph diffusion process introduced in Section 3.3. Below, we describe the initialization of the weight vector $g_{0}$ and the construction of the adjacency matrix $A$.
Initialization of the weight vector.
The initial vector of weights $g_{0}\in\mathbb{R}^{n}$ representing a coarse estimation of the contribution of each Gaussian to the segmentation mask. It is computed by uplifting the 2D foreground mask (either scribbles or a reference mask) from the reference frame into 3D using Eq. (3) and retaining only the top 10% of Gaussians with positive mask values, setting the rest to zero. The nodes for which $g_{0}$ has a positive value define a set of anchor nodes $\mathcal{M}$ that are more likely to contribute to the foreground. The resulting weight vector is a coarse estimation of how much each Gaussian contributes to a rendered 2D segmentation mask.i
Construction of the graph edges.
We define the pairwise similarity function $S_{f}$ as:
@@
where $f_{i},f_{j}$ are the $l_{2}$-normalized DINOv2 features, $s_{f}$ is the median of pairwise $l_{2}$ distances and $b$ is a bandwidth parameter. We choose a global unary regularization term $P(f_{i})$ on each node $i$ contain diffusion to nodes with features similar to those of the foreground. More precisely, $P$ is defined using a similar approach as in Sec. A.2:
@@
The matrix $A$ of graph edges is then defined based on $S_{f}$ and $P$ as in Eq. (7). For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After the $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ iterations). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 00 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding. The selection process of the threshold and bandwidth parameters is detailed below.
Evaluation and hyperparameter tuning.
Segmentation relies on three hyperparameters: the bandwidths for $S_{f}$ and $P$, and the threshold for foreground/background pixel assignment. For SPIn-NeRF, all hyperparameters are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided, hence we predict a SAM mask based on the scribbles of the reference frame, and choose the hyperparameters maximizing the IoU with this SAM mask. This is consistent with a scenario where the user, here SAM, would choose hyperparameters based on visual inspection on one of the frames.

### v5

@@ 5.2 Multi-view segmentation results
Table 1: Multi-view segmentation (IoU) on NVOS [39] and SPIn-NeRF [32].
Geometry only Single view Uplifting Uplifting + Dif.
@@
Table 2: Ablation on SPIn-NeRF segmentation. We compare purely geometrical reference mask reprojection and single-view prediction with our feature/mask uplifting and graph diffusion.
ramen figurines teatime waldo overall
LERF [19] 62.0 75.0 84.8 72.7 73.6
LangSplat [35] 73.2 80.4 88.1 95.5 84.3
LUDVIG (ours) 78.9 80.4 94.9 90.9 86.3
@@
Table 3: LERF object localization. Comparison with the state of the art on the dataset introduced by LangSplat [35].
ramen figurines teatime waldo overall time
Initial evaluation protocol from [19, 35]: 2D object segmentation
LERF [19] 28.2 38.6 45.0 37.9 37.4 45
LangSplat [35] 51.2 44.7 65.1 44.5 51.4 105
LUDVIG (ours) 58.1 63.3 77.1 58.5 64.3 10
New evaluation protocol introduced in OpenGaussian [46]: 3D object selection
OpenGaussian [46] 23.9 59.3 54.4 34.6 43.1 50
Dr. Splat [17] 24.7 53.4 57.2 39.1 43.6 -
LUDVIG (ours) 42.3 58.0 58.6 42.8 50.4 10
Table 4: LERF object segmentation. We evaluate on the LERF dataset introduced by LangSplat [35] with two different evaluation protocols: i) segmentation based on the 2D reprojected features, ii) 3D segmentation (or object selection and reprojection of the binary 3D masks, proposed by [19, 35] and [46] respectively.
5.3 Open-vocabulary object localization
Task. We evaluate on the LERF dataset [19], which includes localization and segmentation tasks on complex in-the-wild scenes. We report our results on the extended evaluation task introduced by LangSplat [35] containing additional challenging localization samples.
Setup. For localization, we simply reproject the 3D CLIP relevancy scores and follow the evaluation protocol from LangSplat [35]. For segmentation, we consider two evaluation protocols: (i) standard 2D segmentation, as introduced by [19] (top part of the table), and (ii) 3D object selection, which consists of binarizing the 3D masks before rendering (bottom part). The latter protocol was introduced in [17] as a way to better assess the quality of 3D semantic features, such as 3D CLIP features, compared to prior existing protocols which focus on the performance on 2D downstream tasks. For both approaches, we refine raw 3D CLIP relevancy scores using graph diffusion and use them for segmentation as follows. For (i) we perform 2D segmentation from rendered features using SAM, as described in Sec. 4.2, which we compare to 2D segmentation by LERF [19] and LangSplat [35] (top of table). For (ii) we directly segment in 3D by binarizing the diffused features before rendering them, then computing 2D segmentation masks from these rendered features. We use automatic thresholding for binarization before and after rendering.
Results. Table 3 reports object localization results, where our method outperforms LangSplat despite not relying on SAM for this task. Table 4 presents segmentation results along with average runtimes, which account for feature map generation and 3D feature training when applicable. Across both evaluation settings, our method consistently surpasses prior approaches while achieving a 5 to 10$\times$ speedup in feature generation and uplifting.
Runtime analysis. Our approach based on graph diffusion offers fast feature map generation compared to leveraging SAM’s automatic mask generation in 2D as in [35, 46], but it comes with additional inference-time overhead. This tradeoff is further discussed in supplementary Sec. B.2. When inference time is a critical constraint, LUDVIG’s uplifting can also be applied directly to full segmentation masks, following the same setup as LangSplat [35] and OpenGaussian [46], as shown in the next section.
@@ 5.3 Open-vocabulary object localization
Table 5: Ablation on LERF object segmentation. Results (IoU) with and without using 3D graph diffusion and/or 2D SAM segmentation, evaluated on the dataset introduced by LangSplat [35].
Ablation study. Table 5 presents an ablation study on the effect of graph diffusion and SAM-based segmentation following the first evaluation protocol (top part) in Table 4.
5.4 Open-vocabulary semantic segmentation
Task. We evaluate on ScanNetv2 [8], which provides posed RGB images from video scans, reconstructed point clouds, and ground-truth 3D point-level semantic labels.
Setup. We follow the evaluation protocol of OpenGaussian [46]. In particular, we train Gaussian Splatting by initializing from the ground-truth point cloud, keeping the 3D positions fixed and disabling densification. For this task, we replace OpenGaussian’s quantized feature learning with our uplifting process, leaving the rest of the protocol unaltered. Specifically, we uplift 2D segmentation masks generated by SAM based on textual queries, as in the prior works we compare to [35, 41, 46].
Results. As shown in Table 6, our approach yields significant accuracy gains over OpenGaussian [46], while being considerably simpler and faster than their quantization-based learning process.
Methods 19 classes 15 classes 10 classes
mIoU mAcc mIoU mAcc mIoU mAcc
LangSplat 3.8 9.1 5.4 13.2 8.4 22.1
LEGaussians 3.8 10.9 9.0 22.2 12.8 28.6
OpenGaussian 24.7 41.5 30.1 48.3 38.3 55.2
LUDVIG 33.9 51.4 37.4 57.2 46.4 66.2
Table 6: ScanNet semantic segmentation. For LUDVIG, we replace OpenGaussian’s quantization-based feature training stage (40$\sim$ 40∼ 40min) with our direct uplifting process (3$\sim$ 3∼ 3min), starting from the same 2D feature maps (50$\sim$ 50∼ 50min) and GS pre-training stage (9$\sim$ 9∼ 9min). Results reported by OpenGaussian.
@@ 6 Concluding remarks and limitations
In this work, we introduce a simple yet effective aggregation mechanism for transferring 2D visual representations into 3D, bypassing traditional optimization-based approaches. The aggregation builds upon already trained Gaussian Splatting representations and is implemented within the CUDA rendering process, making 2D-to-3D uplifting as fast as 3D-to-2D rendering. Our approach can serve as an off-the-shelf module for transferring 2D features into 3D across a wide range of applications, as demonstrated in our experiments (Sec. 5.4) and its recent integration into Panst3R [57] for novel-view panoptic segmentation.
We also propose a graph diffusion process that combines spatial structure with DINOv2 similarity to generate accurate 3D segmentation masks, starting from coarse inputs such as scribbles or CLIP relevancy maps. This leads to strong gains over SAM-based open-vocabulary segmentation baselines while remaining computationally efficient.
However, the quality of the resulting 3D features depends on the underlying 3D scene reconstruction, which remains challenging in cases of high specularity [16, 48] or motion blur [54, 25]. In such scenarios, learning 3D features along with 3D Gaussian Splatting reconstruction may serve as a regularization and improve scene geometry.
@@ Appendix A Using LUDVIG for downstream tasks
In this section, we describe our approach for uplifting DINOv2, SAM and CLIP models and evaluating the 3D features on two downstream tasks: segmentation and open-vocabulary object detection. As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$ and a corresponding 3D scene obtained using the Gaussian Splatting method.
Multi-view segmentation. For this task, we assume that a foreground mask of the object to be segmented is provided on a reference frame taken to be the first frame $I_{1}$. We consider two types of foreground masks: either scribbles or a whole reference mask of the object, both of which define a set of foreground pixels $\mathcal{P}$. In the following subsections, we present the proposed approaches for segmentation using SAM and DINOv2 features, based on both types of foreground masks.
@@ A.1 Multiple-view segmentation with SAM
SAM [21, 37] is a powerful image segmentation model, that can generate object segmentation masks from point prompts on a single 2D image. Aggregating SAM 2D segmentation masks in 3D allows for cross-view consistency and improves single-view segmentation results. In order to leverage SAM, we propose a simple mechanism for generating SAM 2D features for each frame from a foreground mask in the reference frame.
Generating 2D query points for SAM. The key idea is to generate point prompts on each frame from the foreground mask provided on the reference frame. To this end, we perform an uplifting of the foreground mask (Eq. (5) from the main paper) and re-project it on all frames (Eq. (4) from the main paper). From the reprojected mask for viewing direction $d$, further normalized by its average value, we retain a subset $\mathcal{P}_{d}$ of pixels with values higher than a threshold $\tau$ fixed for all scenes ($\tau=0.4$ for SPIn-NeRF and $\tau=1$ for NVOS). We then predict a SAM mask based on these point prompts as described below.
Predicting SAM masks from sets of query points. Given a set of pixels $\mathcal{P}_{d}$ pertaining to the foreground, we compute 2D segmentation masks using SAM by randomly selecting 333 points prompts from $\mathcal{P}_{d}$, repeating the operation 101010 times and averaging the resulting masks for each view to obtain the final 2D SAM feature maps.
Segmentation with uplifted SAM masks. The 2D segmentation masks generated by SAM are uplifted using the aggregation scheme described in Sec. 3.2. Our final prediction is obtained by rendering the uplifted feature maps into the target frame and thresholding.
Evaluation and hyperparameter tuning. Segmentation with 3D SAM masks requires setting a threshold for foreground/background pixel assignment, and optionally choosing one of the three masks proposed by SAM (representing different possible segmentations of the object of interest). For SPIn-NeRF, the threshold and mask prediction are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided; hence, a single mask is predicted, and the segmentation threshold is fixed across all scenes for SAM, and automatically chosen using Li’s iterative Minimum Cross Entropy method [27] for SAM 2.
@@ A.2 Multiple-view segmentation with DINOv2
DINOv2 [33] is a self-supervised vision model recognized for its generalization capabilities. In this work, we aggregate the patch-level representations produced by DINOv2 with registers [9] into a high resolution and fine-grained 3D semantic representation.
Construction of 2D feature maps. We construct the 2D feature maps using a combination of a sliding windows mechanism and dimensionality reduction of the original DINOv2 features. Specifically, we i) extract DINOv2 patch-level representations across multiple overlapping crops of each image, ii) apply dimensionality reduction over the set of all patch embeddings, ii) upsample and aggregate the dimensionality-reduced patch embeddings to obtain pixel-level features for each image. The process is illustrated in Fig. A. This approach enhances the granularity of spatial representations by aggregating patch-level representations to form pixel-level features. To favor the first principal components, known to focus on the foreground objects [33], the features are re-weighted by the eigenvalues of the PCA decomposition.
Segmentation with uplifted DINOv2 features. The 2D feature maps from the $m$ views are uplifted using Eq. (5) from the main paper, and the resulting 3D features are re-projected into any viewing direction $d$ using Eq. (4) from the main paper to compute rendered 2D features $(\hat{F}_{d,p})$. To obtain segmentation masks, we construct a score $P(\hat{F}_{d,p})$ for a 2D pixel $p$ to belong to the foreground, based on its corresponding rendered feature. More precisely, $P$ relies on the rendered foreground features $\mathcal{F}_{\text{ref}}:=(\hat{F}_{d_{1},p})_{p\in\mathcal{P}}$ corresponding to the foreground mask computed on the reference frame $I_{1}$. We propose two approaches for constructing $P$. The first one is a simple approach that sets $P(\hat{F}_{d,p})=\mathcal{S}_{F}(\hat{F}_{d,p},\bar{F})$ where $\bar{F}$ is the average over foreground features $\mathcal{F}_{\text{ref}}$, and $\mathcal{S}_{F}$ is defined based on the cosine similarity. The second approach is more discriminative and first trains a logistic regression model $P$ on all rendered 2D features of the reference frame, so that the foreground features $\mathcal{F}_{\text{ref}}$ are assigned a positive label. Then $P(\hat{F}_{d,p})$ gives the probability that a pixel $p$ belongs to the foreground. The final mask is then obtained by thresholding.
@@
Figure A: Sliding windows for construction of DINOv2 feature maps.
@@ A.3 Enhancing segmentation with DINOv2 using 3D graph diffusion
DINOv2 provides generic visual features that do not explicitly include information for segmentation, unlike models such as SAM that were specifically trained for such a task. Consequently, using the 2D projections of uplifted DINOv2 features, as proposed in Sec. A.2, might fail to separate different objects that happen to have similar features while still being distinct entities. This challenge can be mitigated by incorporating 3D spatial information in which the objects are more likely to be well-separated. To this end, we propose to leverage the graph diffusion process introduced in Sec. 3.3. Below, we describe the initialization of the weight vector $g_{0}$ and the construction of the adjacency matrix $A$.
Initialization of the weight vector. The initial vector of weights $g_{0}\in\mathbb{R}^{n}$ representing a coarse estimation of the contribution of each Gaussian to the segmentation mask. It is computed by uplifting the 2D foreground mask (either scribbles or a reference mask) from the reference frame using Eq. (3) from the main paper, and retaining only the top 10% of Gaussians with positive mask values, setting the rest to zero. The nodes for which $g_{0}$ has a positive value define a set of anchor nodes $\mathcal{M}$ that are more likely to contribute to the foreground. The resulting weight vector is a coarse estimation of how much each Gaussian contributes to a rendered 2D segmentation mask.
Construction of the graph edges. We define the pairwise similarity function $S_{f}$ as:
@@
where $f_{i},f_{j}$ are the $l_{2}$-normalized DINOv2 features, $s_{f}$ is the median of pairwise $l_{2}$ distances and $b$ is a bandwidth parameter. We choose a global unary regularization term $P(f_{i})$ on each node $i$ to contain diffusion to nodes with features similar to those of the foreground. More precisely, $P$ is defined using a similar approach as in Sec. A.2:
@@
The matrix $A$ of graph edges is then defined based on $S_{f}$ and $P$ as in Eq. (7) from the main paper. For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After the $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ iterations). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 0 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding. The selection process of the threshold and bandwidth parameters is detailed below.
Evaluation and hyperparameter tuning. Segmentation relies on three hyperparameters: the bandwidths for $S_{f}$ and $P$, and the threshold for foreground/background pixel assignment. For SPIn-NeRF, all hyperparameters are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided, hence we predict a SAM mask based on the scribbles of the reference frame, and choose the hyperparameters maximizing the IoU with this SAM mask. This is consistent with a scenario where the user, here SAM, would choose hyperparameters based on visual inspection on one of the frames.